OpenStudy (anonymous):

Which of the following could not be points on the unit circle? a.(0.8, -0.6) b.(-2/3,√ 5/3) c.(√ 3/2, 1/3) d.(1,1)

jimthompson5910 (jim_thompson5910):

All points (x,y) on the unit circle satisfy the equation \(\large x^2 + y^2 = 1\)

jimthompson5910 (jim_thompson5910):

The "unit" refers to the radius being r = 1

jimthompson5910 (jim_thompson5910):

Looking at choice A, we see that x = 0.8 and y = -0.6 \[\Large x^2 + y^2 = 1\] \[\Large (0.8)^2 + (-0.6)^2 = 1\] \[\Large 0.64+0.36 = 1\] \[\Large 1 = 1\] So (0.8,-0.6) is definitely on the unit circle

OpenStudy (anonymous):

I was guessing b and c?

jimthompson5910 (jim_thompson5910):

you will do the same basic steps for choices B through D

OpenStudy (anonymous):

I mean A and D

jimthompson5910 (jim_thompson5910):

well I just showed that A is on the unit circle. So choice A is out. They want points that are NOT on the unit circle.

jimthompson5910 (jim_thompson5910):

you need to find points (x,y) that make x^2 + y^2 = 1 false

OpenStudy (anonymous):

Alright got you! thanks

jimthompson5910 (jim_thompson5910):

no problem